OpenStudy (anonymous):

A circle has a diameter of 10 feet. What is the area of the circle? Leave answers in terms of π

OpenStudy (anonymous):

Area of a circle = pi*r^2 \[A = \pi r^2\] Where r is your radius.

OpenStudy (anonymous):

If your diameter is 10 what's your radius?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Good! So how would you put it in your area formula?

OpenStudy (anonymous):

pi*5^2

OpenStudy (anonymous):

Good, so 5^2 =?

OpenStudy (anonymous):

Oh duhh I was thinking of a sphere. X(

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Right so your area = 25pi :)

OpenStudy (anonymous):

\[A = 25\pi\]